# Distributed Algorithms for Potential Problems

### v3

@@ 2.1 Notation and definitions
In this paper, we consider simple undirected graphs unless otherwise stated. For a graph $G=(V,E)$, we denote its set of nodes by $V$ and its set of edges by $E$. If $G$ is not clear from the context we use $V(G)$ to denote the set of nodes of $G$ and $E(G)$ to denote the set of edges of $G$. For any two nodes $u,v\in V(G)$ the distance between $u$ and $v$ is the number of hops in a shortest path between $u$ and $v$, and is denoted by $\operatorname{dist}(u,v)$. For any node $v\in V(G)$, the radius-$r$ neighborhood $\mathcal{N}_{r}[v]$ of $v$ is the set of all nodes at distance at most $r$ from $v$. Let $A\subseteq V(G)$ be a subset of nodes, and $B\subseteq E(G)$ a subset of edges. By $\mathcal{N}_{r}[A]$ we denote the set of all nodes at distance at most $r$ from any node in $A$, and by $\mathcal{N}_{r}[B]$ we denote the set of all nodes that are endpoints of edges in $B$ or are at distance at most $r$ from an endpoint of an edge in $B$. For any two graphs $G,H$, the intersection between $G$ and $H$ is the graph $G\cap H=(V(G)\cap V(H),E(G)\cap E(H))$, and the union between $G$ and $H$ is the graph $G\cup H=(V(G)\cup V(H),E(G)\cup E(H))$. The difference between two graphs $G$ and $H$ is the graph $G\setminus H=(V(G),E(G)\setminus E(H))$.
@@
In this paper, we consider the LOCAL model of distributed computation [Lin92]. We are given a distributed network of $n$ nodes/processors, each of which is capable of arbitrary (but finite) computation. Nodes are connected by edges which represent communication links, and are able to send messages of arbitrarily large (but finite) size to each other: note that nodes can discriminate between different communication links through ports, which are numbered from $1$ to $\Delta$, where $\Delta$ is the maximum degree of the network, that are adversarially assigned. Each node is also given a unique identifier in the set $n^{c}]$ for some constant $c\geq 1$, which is, again, adversarially assigned. The communication network is represented by a simple undirected graph $G=(V,E)$ where $V$ is the set of nodes/processors and $E$ is the set of edges/communication links. All nodes run the same algorithm and computation proceeds in synchronous rounds. In each round, each node sends messages to its neighbors, receives messages from all of its neighbors, and performs local computation. The input of a problem is given in the form of a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of labels. The output of a node is an assignment of labels to the node itself and/or to its incident half-edges: note that this information can be stored in the node through the use of port numbers. When a node determines its own output, we say that the node terminates its computation. The running time of a LOCAL algorithm is the number of communication rounds it takes for all nodes to produce their output. We also refer to the running time of an algorithm as the locality of the algorithm. The complexity of a problem is the minimum running time (across all possible algorithms) needed to solve the problem. We can also assume that nodes are equipped with some source of randomness (independently of other nodes), such as an arbitrarily long random bit string, which they can use during the computation. In this case, we talk about the randomized LOCAL model as opposed to the deterministic LOCAL model. In case of randomized computation, we require that the problem is solved with probability at least 11/n1-1/n.
@@ 2.2 Problems of interest
Definition 2.5 (Labeled graph satysfying a set of constraints).
@@ A Locally Optimal Problem (LOP) is a pair $(\Pi,\Psi)$ where:
Moreover, $\mathcal{C}$ must satisfy the following property, which in essence says that a correct solution does not allow local improvements. For any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$, let us denote by $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ the labeling of $G$. We say that the potential associated to $(G,\ell)$ is $\operatorname{Pot}(G,\ell)=\sum_{v\in V}\Psi_{\ell}(G_{r}(v),v)$, where $\Psi_{\ell}(G_{r}(v),v)$ is the evaluation of $\Psi$ on $(G_{r}(v),v)$ that is labeled by $\ell$. For any node $v\in V$, let the cost of $v$ be $c_{v}:=\Psi_{\ell}(G_{r}(v),v)$. Then, the centered graph $(G_{2r}(v),v)$ labeled with $\ell$ belongs to $\mathcal{C}$ if and only if we cannot change the labeling of $v$ to improve the potential of a node in $(G_{2r}(v),v)$. Formally, there is no labeling $\ell^{\prime}$ with the following properties, where, for any node $u\in V$, $c_{u}^{\prime}:=\Psi_{\ell^{\prime}}(G_{r}(u),u)$.
  • $\ell^{\prime}$ differs from $\ell$ only on the output labels of $v$ and of its incident half-edges.
  • According to the labeling $\ell^{\prime}$, $G_{2r}(v)$ is still a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph.
@@ A Generalized Locally Optimal Problem (GLOP) is a pair $(\Pi,\Psi)$ where:
  • The centered graph $(G_{r}(v),v)$ labeled with $\ell^{\prime}$ belongs to $\mathcal{C}$.
@@
In other words, if a centered graph $(G_{3r}(v),v)$ labeled with $\ell$ is not contained in $\mathcal{C}$, then the outputs on $(G_{r}(v),v)$ can be changed such that we obtain a centered graph in $\mathcal{C}$, and the obtained potential is strictly better. (Observe that the changes on $G_{r}(v)$ can only affect nodes within distance $2r$ from $v$, and their neighborhood is contained in the radius-$3r$ neighborhood of $v$.)
@@
  • The sets of input labels is the same, i.e., $\mathcal{V}_{\operatorname{in}}^{\prime}=\mathcal{V}_{\operatorname{in}}$, $\mathcal{E}_{\operatorname{in}}^{\prime}=\mathcal{E}_{\operatorname{in}}$.
@@
  • The radius of the centered graphs in $\mathcal{C}^{\prime}$ is $3r^{\prime}=6r$ and the maximum degree is $\Delta^{\prime}=\Delta$.
@@
Let $(p_{1},X_{1},Y_{1}),\ldots,(p_{\deg(v)},X_{\deg(v)},Y_{\deg(v)})$ be the labels assigned by $\ell$ to the half-edges incident to $v$. The values $p_{1},\ldots,p_{\deg(v)}$ are called ports of $v$. If $\exists i,j~:~X_{i}\neq X_{j}\text{ or }Y_{i}\neq Y_{j}$, then $f((G_{r}(v),v))$ is undefined. Let $X=X_{1}$ and $Y=Y_{1}$. Consider a BFS visit on $G_{r}(v)$ where, the order in which the neighbors of some node $u$ are visited, is according to the ports of $u$. If $u$ has ports with values not in $\deg(u)]$, or two incident ports with the same value, then $f$ is undefined. The labeling $f$ on $(G_{r}(v),v)$ is then obtained by assigning, to the $j$-th visited half-edge, the $j$-th element of $X$, and to the $j$-th visited node, the $j$-th element of $Y$.
We now define a function $g$ that takes as input a centered graph $(G_{2r}(v),v)$ labeled with $\ell$ and returns a labeling for $(G_{r}(v),v)$ that uses the labels of $\Pi$. By abusing the notation, with $g$ we also denote the obtained labeling. Again, the function $g$ will be a partial function. For each node $u$ in $(G_{r}(v),v)$, compute $f((G_{r}(u),u))$. If $f$ is undefined in any of the cases in which we applied it, then $g$ is undefined on $(G_{2r}(v),v)$. For each node (resp. half-edge) $z$ in $(G_{r}(v),v)$, the labeling $g$ for $\Pi$ is defined as follows. Let $a$ be the array containing the output that each node in $(G_{r}(v),v)$ produced for $z$, according to $f$. Then, $g$ labels $z$ with the $(\sum_{x\in a}x\pmod{|\mathcal{V}_{\operatorname{out}}|})-$th label in $\mathcal{V}_{\operatorname{out}}$ (resp. $(\sum_{x\in a}x\pmod{|\mathcal{E}_{\operatorname{out}}|})-$th label in $\mathcal{E}_{\operatorname{out}}$), according to some arbitrary but fixed order.
@@
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=\Psi_{g}(G_{r}(v),v)\text{ if $g$ on $(G_{r^{\prime}=2r}(v),v)$ is defined, and}$
$\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v):=M\text{ otherwise,}$
@@
where $M$ is an arbitrary value strictly larger than any value returned by $\Psi$.
@@
The constraint $\mathcal{C}^{\prime}$ contains exactly those $\ell$-labeled centered graphs $(G_{3r^{\prime}}(v),v)$ such that there is no labeling $\ell^{\prime}$, that differs only on the output labels of $v$ and its incident half-edges, that satisfies $\Psi^{\prime}_{\ell^{\prime}}(G_{r^{\prime}}(v),v)<\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v)$ and $\operatorname{Pot}(G_{r^{\prime}}(v),\ell^{\prime})<\operatorname{Pot}(G_{r^{\prime}}(v),\ell)$. This condition ensures that the LOP $(\Pi^{\prime},\Psi^{\prime})$ that we defined satisfies the requirements of Definition 2.7.
Now, we prove that any solution for $(\Pi^{\prime},\Psi^{\prime})$ can be converted into a solution for $(\Pi,\Psi)$ in $O(1)$ rounds. We first observe that, in any valid solution for $(\Pi^{\prime},\Psi^{\prime})$, the potential on each node $v$ must be strictly less than $M$. In fact, if the potential computed on node $v$ is $M$, it means that $g$ is undefined on the $r^{\prime}=2r$-radius neighborhood of $v$. This implies that $f$ is undefined on some node $u$ within distance $r$ from $v$, and the potential on $u$ is hence also $M$. Observe that, in this case, $u$ can improve its potential by producing an output that encodes a valid port assignment of its $r$-hop neighborhood. ∎

### v4

@@ 2.1 Notation and definitions
In this paper, we consider simple undirected graphs unless otherwise stated. For a graph $G=(V,E)$, we denote its set of nodes by $V$ and its set of edges by $E$. If $G$ is not clear from the context we use $V(G)$ to denote the set of nodes of $G$ and $E(G)$ to denote the set of edges of $G$. For any two nodes $u,v\in V(G)$ the distance between $u$ and $v$ is the number of hops in a shortest path between $u$ and $v$, and is denoted by $\operatorname{dist}(u,v)$. For any node $v\in V(G)$, the radius-$r$ neighborhood $\mathcal{N}_{r}[v]$ of $v$ is the set of all nodes at distance at most $r$ from $v$. Let $A\subseteq V(G)$ be a subset of nodes, and $B\subseteq E(G)$ a subset of edges. By $\mathcal{N}_{r}[A]$ we denote the set of all nodes at distance at most $r$ from any node in $A$, and by $\mathcal{N}_{r}[B]$ we denote the set of all nodes that are endpoints of edges in $B$ or are at distance at most $r$ from an endpoint of an edge in $B$. Similarly, by $\mathcal{N}_{r}(v)$ we denote the radius-$r$ (open) neighborhood of $v$, which is defined simply as $\mathcal{N}_{r}(v)=\mathcal{N}_{r}[v]\setminus\{v\}$. When we omit the subscript $r$, we assume that $r=1$. For any two graphs $G,H$, the intersection between $G$ and $H$ is the graph $G\cap H=(V(G)\cap V(H),E(G)\cap E(H))$, and the union between $G$ and $H$ is the graph $G\cup H=(V(G)\cup V(H),E(G)\cup E(H))$. The difference between two graphs $G$ and $H$ is the graph $G\setminus H=(V(G),E(G)\setminus E(H))$.
@@
In this paper, we consider the LOCAL model of distributed computation [39]. We are given a distributed network of $n$ nodes/processors, each of which is capable of arbitrary (but finite) computation. Nodes are connected by edges which represent communication links, and are able to send messages of arbitrarily large (but finite) size to each other: note that nodes can discriminate between different communication links through ports, which are numbered from $1$ to $\Delta$, where $\Delta$ is the maximum degree of the network, that are adversarially assigned. Each node is also given a unique identifier in the set $n^{c}]$ for some constant $c\geq 1$, which is, again, adversarially assigned. The communication network is represented by a simple undirected graph $G=(V,E)$ where $V$ is the set of nodes/processors and $E$ is the set of edges/communication links. All nodes run the same algorithm and computation proceeds in synchronous rounds. In each round, each node sends messages to its neighbors, receives messages from all of its neighbors, and performs local computation. The input of a problem is given in the form of a $(\mathcal{V},\mathcal{E})$-labeled graph $G=(V,E)$ where $\mathcal{V}$ and $\mathcal{E}$ are sets of labels. The output of a node is an assignment of labels to the node itself and/or to its incident half-edges: note that this information can be stored in the node through the use of port numbers. When a node determines its own output, we say that the node terminates its computation. The running time of a LOCAL algorithm is the number of communication rounds it takes for all nodes to produce their output. We also refer to the running time of an algorithm as the locality of the algorithm. The complexity of a problem is the minimum running time (across all possible algorithms) needed to solve the problem. We can also assume that nodes are equipped with some source of randomness (independently of other nodes), such as an arbitrarily long random bit string, which they can use during the computation. In this case, we talk about the randomized LOCAL model as opposed to the deterministic LOCAL model. In case of randomized computation, we require that the problem is solved with high probability (w.h.p.), that is, with probability at least 11/n1-1/n. We remark that in case of randomized LOCAL, we assume that unique identifiers are generated by the nodes uniformly at random. This can be done with probability $1-1/n^{c}$ for arbitrarily large constant $c$ by generating identifiers of length $O(\operatorname{log}n)$ bits, and thus does not affect the success probability of an algorithm.
@@ 2.2 Problems of interest
Definition 2.5 (Labeled graph satisfying a set of constraints).
@@ A Locally Optimal Problem (LOP) is a pair $(\Pi,\Psi)$ where:
Moreover, $\mathcal{C}$ must satisfy the following property, which in essence says that a correct solution does not allow local improvements. For any $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph $G=(V,E)$, let us denote by $\ell=(\ell_{\operatorname{nodes}},\ell_{\operatorname{half-edges}})$ the labeling of $G$. We say that the potential associated to $(G,\ell)$ is $\operatorname{Pot}(G,\ell)=\sum_{v\in V}\Psi_{\ell}(G_{r}(v),v)$, where $\Psi_{\ell}(G_{r}(v),v)$ is the evaluation of $\Psi$ on $(G_{r}(v),v)$ that is labeled by $\ell$. For any node $v\in V$, let the cost of $v$ be $c_{v}:=\Psi_{\ell}(G_{r}(v),v)$. Then, the centered graph $(G_{2r}(v),v)$ labeled with $\ell$ belongs to $\mathcal{C}$ if and only if we cannot change the labeling of $v$ to improve the potential of a node in $(G_{2r}(v),v)$. Formally, there is no labeling $\ell^{\prime}$ with the following properties, where, for any node $u\in V$, $c_{u}^{\prime}:=\Psi_{\ell^{\prime}}(G_{r}(u),u)$:
  • Labeling $\ell^{\prime}$ differs from $\ell$ only on the output labels of $v$ and of its incident half-edges.
  • According to labeling $\ell^{\prime}$, $G_{2r}(v)$ is still a $(\mathcal{V}_{\operatorname{in}}\times\mathcal{V}_{\operatorname{out}},\mathcal{E}_{\operatorname{in}}\times\mathcal{E}_{\operatorname{out}})$-labeled graph.
@@ A Generalized Locally Optimal Problem (GLOP) is a pair $(\Pi,\Psi)$ where:
  • The centered graph $(G_{3r}(v),v)$ labeled with $\ell^{\prime}$ belongs to $\mathcal{C}$.
  • Let $c_{v}:=\Psi_{\ell}(G_{r}(v),v)$ and $c^{\prime}_{v}:=\Psi_{\ell^{\prime}}(G_{r}(v),v)$. It holds that $c^{\prime}_{v}<c_{v}$.
@@
In other words, if a centered graph $(G_{3r}(v),v)$ labeled with $\ell$ is not contained in $\mathcal{C}$, then the outputs on $(G_{r}(v),v)$ can be changed such that we obtain a centered graph in $\mathcal{C}$, and the obtained potential is strictly better, globally and at node $v$. (Observe that the changes on $G_{r}(v)$ can only affect nodes within distance $2r$ from $v$, and their neighborhood is contained in the radius-$3r$ neighborhood of $v$.)
@@
  • The sets of input labels are the same, i.e., $\mathcal{V}_{\operatorname{in}}^{\prime}=\mathcal{V}_{\operatorname{in}}$, $\mathcal{E}_{\operatorname{in}}^{\prime}=\mathcal{E}_{\operatorname{in}}$.
@@
  • Let $r^{\prime}:=2r$. The radius of the centered graphs in $\mathcal{C}^{\prime}$ is $2r^{\prime}=4r$ and the maximum degree is $\Delta^{\prime}=\Delta$.
@@
Let $(p_{1},X_{1},Y_{1}),\ldots,(p_{\deg(v)},X_{\deg(v)},Y_{\deg(v)})$ be the labels assigned by $\ell$ to the half-edges incident to $v$. The values $p_{1},\ldots,p_{\deg(v)}$ are called ports of $v$. If $\exists i,j~:~X_{i}\neq X_{j}\text{ or }Y_{i}\neq Y_{j}$, then $f((G_{r}(v),v))$ is undefined. Let $X=X_{1}$ and $Y=Y_{1}$. Consider a BFS visit on $G_{r}(v)$ where, the order in which the neighbors of some node $u$ are visited, is according to the ports given as output by $u$. If $u$ has ports with values not in $\deg(u)]$, or two incident ports with the same value, then $f$ is undefined. The labeling $f$ on $(G_{r}(v),v)$ is then obtained by assigning, to the $j$-th visited half-edge, the $j$-th element of $Y$, and to the $j$-th visited node, the $j$-th element of $X$. For each node $v$, we say that $f$ succeeds at $v$ if $f((G_{r}(v),v))$ is defined.
We now define a function $g$ that takes as input a centered graph $(G_{2r}(v),v)$ labeled with $\ell$ and returns a labeling for $(G_{r}(v),v)$ that uses the labels of $\Pi$. By abusing the notation, with $g$ we also denote the obtained labeling. Again, the function $g$ will be a partial function. For each node $u$ in $(G_{r}(v),v)$, compute $f((G_{r}(u),u))$. If $f$ is undefined in any of the cases in which we applied it, then $g$ is undefined on $(G_{2r}(v),v)$. For each node (resp. half-edge) $z$ in $(G_{r}(v),v)$, the labeling $g$ for $\Pi$ is defined as follows. Let $a$ be the array containing the output that each node in $(G_{r}(v),v)$ produced for $z$, according to $f$. Then, $g$ labels $z$ with the $(\sum_{x\in a}x\pmod{|\mathcal{V}_{\operatorname{out}}|})-$th label in $\mathcal{V}_{\operatorname{out}}$ (resp. $(\sum_{x\in a}x\pmod{|\mathcal{E}_{\operatorname{out}}|})-$th label in $\mathcal{E}_{\operatorname{out}}$), according to some arbitrary but fixed order. For each node $v$, we say that $g$ succeeds at $v$ if $g((G_{2r}(v),v))$ is defined.
@@
[…]
@@
where $M^{\prime}$ and $M$ are arbitrary values strictly larger than any value returned by $\Psi$, and such that $M^{\prime}>M$.
@@
The constraint $\mathcal{C}^{\prime}$ contains exactly those $\ell$-labeled centered graphs $(G_{2r^{\prime}}(v),v)$ such that there is no labeling $\ell^{\prime}$, that differs only on the output labels of $v$ and its incident half-edges, that satisfies $\Psi^{\prime}_{\ell^{\prime}}(G_{r^{\prime}}(v),v)<\Psi^{\prime}_{\ell}(G_{r^{\prime}}(v),v)$ and $\operatorname{Pot}(G_{r^{\prime}}(v),\ell^{\prime})<\operatorname{Pot}(G_{r^{\prime}}(v),\ell)$. This condition ensures that the LOP $(\Pi^{\prime},\Psi^{\prime})$ that we defined satisfies the requirements of Definition 2.7.
Now, we prove that any solution for $(\Pi^{\prime},\Psi^{\prime})$ can be converted into a solution for $(\Pi,\Psi)$ in $O(1)$ rounds.
We first observe that, in any valid solution for $(\Pi^{\prime},\Psi^{\prime})$, the potential on each node $v$ must be strictly less than $M^{\prime}$. Assume for a contradiction that some node has potential $M^{\prime}$. Then there exists some node $v$ with some local issue (either an invalid port assignment, or inconsistent arrays assigned to the half-edges of $v$), and $f$ does not succeed at $v$. Consider the following improvement:
  • If the ports at $v$ are invalid, $v$ considers some arbitrary order of its incident half-edges, and outputs, for each $1\leq i\leq\deg(v)$, $(i,X,Y)$ on its $i$th half-edge, using the same arbitrary arrays $X,Y$ on all incident half-edges.
  • If the ports at $v$ are valid, for each $1\leq i\leq\deg(v)$, let $e_{i}$ be the incident half-edge that currently has port value $i$. Node $v$ outputs $(i,X,Y)$ on $e_{i}$, using the same arbitrary arrays $X,Y$ on all incident half-edges. Note that, with this change, the port assignment remains unchanged.
Observe that, in both cases, the potential of $v$, after this change, drops from $M^{\prime}$ to at most $M$. Moreover, since the ports on $v$ are not changed if already correct, we get that, if before the change $f$ succeeded at some node $u$ with $\operatorname{dist}(u,v)\leq r$, then after the change $f$ still succeeds at $u$ (and its evaluation is unchanged). Hence, if the ports of $v$ were valid before the change, then nodes in $G_{r^{\prime}}(v)\setminus G_{r}(v)$ are unaffected by the change, while if the ports of $v$ were invalid, then their potential can only stay the same or improve. Moreover, for all nodes within distance $r$ from $v$, $g$ was undefined before the change (since $f$ was undefined at $v$), and hence had potential at least $M$. After the change, each such node has potential no larger than before. Hence, no node worsens its potential, while $v$ improves from $M^{\prime}$ to at most $M$, a contradiction.
We now prove that, in any valid solution for $(\Pi^{\prime},\Psi^{\prime})$, the potential on each node $v$ must be strictly less than $M$. We already know that the potential on each node is at most $M$. Suppose for a contradiction that there is some node $v$ with potential exactly $M$. Then, $f$ succeeds at $v$, but $g$ does not. Hence, there exists some node $u$, within distance $r$ from $v$, where $f$ does not succeed. This would imply that $u$ has potential $M^{\prime}>M$, a contradiction.
The algorithm that converts a solution for $(\Pi^{\prime},\Psi^{\prime})$ into a solution for $(\Pi,\Psi)$ is defined as follows. Each node $v$ gathers its $2r$-hop neighborhood and computes $g((G_{2r}(v),v))$, and outputs the label that $g$ assigns to $v$ and its incident half-edges. This algorithm requires $O(1)$ rounds, since $r$ is constant. Let $\ell^{\prime}$ be the labeling of the given solution for $(\Pi^{\prime},\Psi^{\prime})$, and let $\ell$ be the labeling for $\Pi$ obtained by decoding $\ell^{\prime}$ with $g$. Assume for a contradiction that $\ell$ is not a solution for $\Pi$. Then there exists some node $v$ satisfying that $(G_{3r}(v),v)\notin\mathcal{C}$. By the definition of GLOP, this means that there exists a relabeling $\hat{\ell}$ of $G_{r}(v)$ such that, according to $\hat{\ell}$, $(G_{3r}(v),v)\in\mathcal{C}$, $c^{\prime}_{v}<c_{v}$ (where $c_{v}:=\Psi_{\ell}(G_{r}(v),v)$ and $c^{\prime}_{v}:=\Psi_{\hat{\ell}}(G_{r}(v),v)$), and $\operatorname{Pot}(G_{2r}(v),\hat{\ell})<\operatorname{Pot}(G_{2r}(v),\ell)$. Now modify $\ell^{\prime}$ as follows: change only the output arrays $(X,Y)$ on node $v$, such that $g$ produces exactly the labels of $\hat{\ell}$ on $G_{r}(v)$ (this is possible, since the output of $g$ is a modular sum). Let $\ell^{\prime\prime}$ be the obtained labeling. Since all nodes have potential strictly smaller than $M$, $g$ succeeds at every node. Therefore $\Psi^{\prime}_{\ell^{\prime\prime}}(G_{r^{\prime}}(v),v)=c^{\prime}_{v}<c_{v}=\Psi^{\prime}_{\ell^{\prime}}(G_{r^{\prime}}(v),v)$, and, because only nodes in $G_{r^{\prime}}(v)$ are affected and the labeling decoded by $g$ is exactly $\hat{\ell}$ on $G_{r}(v)$, $\operatorname{Pot}(G_{r^{\prime}}(v),\ell^{\prime\prime})<\operatorname{Pot}(G_{r^{\prime}}(v),\ell^{\prime})$. Hence $v$ would have a valid local improving move in $\Pi^{\prime}$, contradicting that $\ell^{\prime}$ is a solution to $\Pi^{\prime}$. ∎
The definition of locally checkable labeling problems inherently asks for the maximum degree $\Delta$ of the input graph to be a constant. However, note that all definitions in this section generalize to the case where $\Delta=\Delta(n)$ is a function of $n$. In our analysis, we will consider the general case where $\Delta$ is indeed a function of $n$.